OpenStudy (anonymous):

18. Solve the system of equations 4x – y = 9 and 3x – 5y = 11. What is the x-coordinate of the solution? A. –2 B. –1 C. 2 D. 3

OpenStudy (anonymous):

4x-y= 9 ... 1 3x-5y=11 ... 2 2*5: 20x-5y=45 ...2' 2'-1: 17x=34 x=2 Therefore, the x-coordinate of the solution is 2. Also, 8-y=9, y=-1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome(: